OpenStudy (anonymous):

Use an equation to solve. 8 is what percent of 2?

OpenStudy (unanimoose):

^ No, that is wrong.

OpenStudy (unanimoose):

The question is "8 is what precent of 2", not "8 percent of 2."

myininaya (myininaya):

replace the is word with = replace the of word with * to get an equation to solve like 8 is what percent of 2 8 = n * 2 8=n*2 or 8=2n or you can try to remember to solve these equations by doing: n is what percent of m? always do n/m and you can write as a percent after that

OpenStudy (wmj259):

How many times does 2 go into 8?

OpenStudy (anonymous):

4

OpenStudy (shaadykhamlichi):

sorry it's 400

OpenStudy (shaadykhamlichi):

why r u reporting me and at the same time y r giving answers

OpenStudy (wmj259):

Now because 8 is greater then 2 it must be more then 100% correct?

OpenStudy (wmj259):

CoC wants you to not give out direct answers.

OpenStudy (uri):

 convert english to math